When I was 4 - my brother was half my age,
now I am 50,
how old is he?

Answer by Maths68(1474)

You were = 4 years
Your brother was = 2 years (half of your age)
You are = (4+46) =50 years old now
Your brother is = (2+46) = 48 years old now